OpenStudy (anonymous):

if=3/7, q=1/5 and r=2/5 then 14p-9r?^2/4q^2=

OpenStudy (anonymous):

Go ahead and substitute them into your equation, as such. p = 3/7 q = 1/5 r = 2/5

OpenStudy (anonymous):